[Hallc_running] [New Logentry] RC daily update

simona at jlab.org simona at jlab.org
Thu May 3 12:40:04 EDT 2018

Logentry Text:
About 20 minutes before 5 pm yesterday the main router went down and this impacted operations severely. 

Here is a list of entries that explain the symptoms and consequences of what happened on the accelerator and cryo side and show the progression of the recovery so far:

general entry from accelerator operator summarizing what happened: https://logbooks.jlab.org/entry/3571073

updates on the router recovery: https://logbooks.jlab.org/entry/3571121 and  https://logbooks.jlab.org/entry/3571124

update on controls recovery: https://logbooks.jlab.org/entry/3571212 , https://logbooks.jlab.org/entry/3571273

updates on cryo: https://logbooks.jlab.org/entry/3571068, https://logbooks.jlab.org/entry/3571058

On the Hall C side and related to the router outage:  we lost target controls for a while but the targets stayed stable (we switched loop 2 and 3 to the aux HPH). The HMS and SHMS magnets suffered no impact from the incident.

Unrelated to the router issue, cdaql3 hung few minutes before 4 pm and it had to be rebooted (the reboot was completed in about 2 hours).

>From MCC meeting this morning:

Recovery from the router outage continues. The accelerator experienced some hardware failures and those were being addressed. The plan was to continue with the pass change for Hall A and go through beam restoration to all halls  once all the issues are addressed (a list of issues to be addressed from the PD is shown here: https://accweb.acc.jlab.org/presenter/presentation/5523#2 ) . 
The best guess for beam restoration to Hall C as of this morning was sometime during the swing shift. 

Since the MCC 8 am meeting the CHL1 tripped (around 9 am) and now the South Linac is down. I just got an update from the Crew Chief: the pump-down of the SL started about 1 hour ago. The hope is still to restore beam to C sometime during the swing shift.

Lastly, there was more work done on the SHMS Heavy Gas Cherenkov yesterday during the controlled access: the bottom PMTs were replaced with a quartz window Photonis PMT and a borosilicate window PMT with wavelength shifter. The PMTs have seen no beam yet (in C).

Hall C Daily RC meeting: 1:30 p.m. in the conference room located on the second floor of the counting house.

Remote participation via Blue Jeans:
browser: https://bluejeans.com/543903672

Meeting ID: 543903672
phone: (888) 240-2560 (US toll free)

list of international numbers: https://bluejeans.com/numbers



This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3571358
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180503/2e6c92f4/attachment.html>

More information about the Hallc_running mailing list